OpenStudy (anonymous):

"Write each of the following in terms of sin(theta) and cos (theta) then simplify if possible. ..." Sin(theta)cot(theta)+cos(theta)

OpenStudy (john_es):

Do you mean, \[\sin\theta\cos\theta+\cos\theta=\cos\theta(\sin\theta+1)\]

OpenStudy (anonymous):

No, my worksheet says \[\sin \theta \cot \theta + \cos \theta \]

OpenStudy (john_es):

\[\sin\theta\cot\theta+\cos\theta=\sin\theta\frac{\cos\theta}{\sin\theta}+\cos\theta=2\cos\theta\]

OpenStudy (anonymous):

Yes but how did get to cos(theta) ? I need to look at it step by step?

OpenStudy (john_es):

Well, the definition of \[\cot\theta\]is \[\cot\theta=\frac{1}{\tan\theta}\frac{1}{\sin\theta/\cos\theta}=\frac{\cos\theta}{\sin\theta}\]

OpenStudy (john_es):

There must be an equal between the second and the third fraction I wrote. Sorry.

OpenStudy (anonymous):

I worked mine different :/

OpenStudy (anonymous):